OpenStudy (anonymous):

Adding Rational Expressions. (WITH FACTORING) [Medal For Best Descriptive Answer♥]

OpenStudy (ivancsc1996):

Which is the question?

OpenStudy (ivancsc1996):

Which is the denominator of the second term?

OpenStudy (ivancsc1996):

x-5

OpenStudy (anonymous):

pfft sorry about that.

OpenStudy (anonymous):

\[\frac{ 15-8x }{ x ^{2} -5x} + \frac{ x }{ x-5 }\]

OpenStudy (ivancsc1996):

Well, this is what you need to do\[\frac{ (15-8x)(x-5) }{ (x ^{2} -5x)(x-5)}+\frac{ x(x ^{2}-5x) }{ (x-5)(x ^{2} -5x) }=\frac{ (15-8x)(x-5)+x(x ^{2}-5x) }{ (x ^{2} -5x)(x-5) }\]\[\frac{ 15x-75-8x ^{2}+40x+x ^{3}-5x ^{2} }{ x ^{3}-5x ^{2}-5x ^{2}+25x }=\frac{ x ^{3}-13x ^{2}+55x-75 }{ x ^{3}-10x ^{2}+25x }\]

OpenStudy (anonymous):

o_o

OpenStudy (anonymous):